A baseball team has home games on Friday and saturday. The two games together earn $ 4036.00 for the team.?
Friday's game generates $1124.00 less than saturday's game. How much money was taken in at each game?

Answer by Maths68(1474)

Earns on Friday = f
Earns on Saturday = s
Total Earning = $4036
Friday's game generates $1124.00 less than saturday's game
f=s-1124........(1)
Total Earning = Earns on Friday + Earns on Saturday
4036=f+s
Put the value of s from (1) to above equation
4036=f+s
4036=(s-1124)+s
4036=s-1124+s
4036=2s-1124
4036+1124=2s
5160=2s
Divide by 2 both sides of the equation
5160/2=2s/2
2580=s
s=2580
Put the value of s in (1)
f=s-1124........(1)
f=2580-1124
f=1456
Earns on Friday = f = 1456
Earns on Saturday = s = 2580